APPLIED MATHEMATICS 2130

Introductory Numerical Techniques

Fall 1999

Slot Section Classroom  Instructor (Office)
17  001  HH-3030 Dr. S. Kocabiyik (HH-3032)

Prerequisites

M 1001 and AM 2120 (or CS 2710 or CS 2602)

Text

There is an “Applied Mathematics 2130 Manual” for sale in the bookstore.

Evaluation

Grades will be based on four reports (assignments). Both the information contained in the reports and the quality of exposition will be taken into account in the evaluation. There will be no final examination. The values of the reports are 15 marks, 25 marks, 30 marks and 30 marks for reports 1–4, respectively. A schedule for submission of lab reports is presented below. There is a penalty associated with late submissions resulting in a deduction of 5 marks per week (or any fraction thereof) of lateness. Consult the first section of the course manual for additional details.

A main purpose of the course is to teach technical writing. After each report has been read, you will meet with your instructor to go over your submission. A report is not assigned any grade until after this meeting has taken place. At this meeting, students must be prepared to explain the workings of any computer program submitted, and such explanations, if requested, will contribute to the grade on the project.

Lab Report Schedule

Report Due
1  9:00 a.m. Tuesday, October 5
2  9:00 a.m. Tuesday, October 26
3  9:00 a.m. Tuesday, November 16
4  9:00 a.m. Monday, December 6

Office Hours

Tuesdays and Thursdays from 11:00 a.m. to 12:30 p.m. - I am also available at other times by appointment.

Web address

A web page with course information, handouts, and extras is regularly maintained. The address is

                                 http://www.math.mun.ca/^{∼}m2130/.